# Independent connection in ACTion during inflation

### v1

@@
Independent connection in ACTion during inflation
Alberto Salvio
Physics Department, University of Rome and INFN Tor Vergata, Italy
The Atacama Cosmology Telescope (ATC) has recently released new measurements and constraints on inflationary observables. In this paper it is shown that a component of a dynamical affine connection, which is independent of the metric, can easily drive inflation in agreement with these observations. Such geometrical explanation of inflation is analysed in detail here in the minimal model, including the predictions for the scalar spectral index $n_{s}$ and its running $\alpha_{s}$, the amplitude of the scalar perturbations and the tensor-to-scalar ratio $r$. Furthermore, this minimal model is shown to provide an inflationary attractor: arbitrary initial values of the kinetic energy density are dynamically attracted down to negligible values compared to the potential energy density in homogeneous and isotropic metrics. The role of the Higgs boson during and after inflation is also briefly discussed.
@@ I Introduction
In Refs. [1, 2] the ATC collaboration has provided new determinations of several inflationary observables. One of the most evident differences with respect to the values previously furnished by the Planck and BICEP/Keck (BK18) collaborations [3, 4] is a higher favored value for $n_{s}$: when the new ATC observations are combined with Planck data and baryon acoustic oscillation data from the Dark Energy Spectroscopic Instrument (DESI) [5] (this combination is denoted P-ACT-LB) Ref. [1] finds $n_{s}=0.9743\pm 0.0034$.
@@
The purpose of this paper is to offer a well-motivated scenario that can explain the new results of Refs. [1, 2]. While Einstein’s general relativity (GR) explains gravity in geometrical terms (the distances are measured through the metric and the gravitational force is determined by the affine connection), inflation is usually accounted for by adding an ad hoc scalar field, which is unrelated to geometry. However, from the purely geometrical point of view the metric and the connection, unlike in GR, can be completely independent objects and, moreover, can contain extra degrees of freedom besides the spin-2 graviton. This generalized scenario is known as metric-affine gravity (see [12, 13] for detailed discussions and further references). The extra degrees of freedom carried by the affine connection can include the inflaton [14, 15, 16] and this provides a geometrical origin also for this scalar field and the subsequent period of reheating too [14, 15], which sets the initial conditions for the radiation dominated phase of the universe.
@@ II The model
is the curvature built with the affine connection ${\cal A}_{\mu\leavevmode\nobreak\ \sigma}^{\leavevmode\nobreak\ \,\rho}$, also $\epsilon^{\mu\nu\rho\sigma}$ is the totally antisymmetric Levi-Civita symbol with $\epsilon^{0123}=1$ and $g$ is the determinant of the metric. The invariants $\mathcal{R}$ and $\mathcal{R}^{\prime}$ generically appear in low-energy metric-affine effective field theories [21]. In the GR case, where ${\cal A}_{\mu\leavevmode\nobreak\ \sigma}^{\leavevmode\nobreak\ \,\rho}$ equals the Levi-Civita connection, ${\cal R}$ coincides with the Ricci scalar, $R$, but ${\cal R^{\prime}}$ vanishes. Thus in metric-affine gravity ${\cal R^{\prime}}$ can be understood as a component of the connection. The coefficients $\alpha$ and $\beta$ have dimensions of energy squared, while $c$ is a dimensionless coefficient. The quantity $M_{P}^{2}/(4\beta)$ is called the Barbero-Immirzi parameter [22, 23]. These coefficients can be determined as follows.
@@
Despite its simplicity, this minimal model predicts inflationary observables that are quite robust under extensions of the inflationary action in (1). First, the same inflationary predictions emerge if one substitutes $c{\cal R^{\prime}}^{2}$ with a general quadratic function of both ${\cal R}$ and ${\cal R^{\prime}}$. This is because such more general function leads to the same potential [13] (see also [26]). Second, a variation of this idea where the action includes only a quadratic function of both ${\cal R}$ and ${\cal R^{\prime}}$, and so it is invariant under $g_{\mu\nu}\to\Omega^{2}g_{\mu\nu}$, for a generically spacetime dependent $\Omega$ (Weyl invariance), leads again to the same potential [16].
@@ III Inflationary predictions
[…]
@@
$\xi_{2}(\omega)=\frac{64M_{P}^{4}\cosh^{2}(X(\omega))\left(M_{P}^{2}\sinh(X(\omega))-\beta\right)}{9\left(M_{P}^{2}\sinh(X(\omega))-4\beta\right)^{3}}.\$ (13)
@@
$\begin{array}[]{c}\left\{\begin{array}[]{c}\Pi^{\prime}=-3\Pi-\frac{1}{H}\frac{d\mathcal{U}}{d\omega},\\ \\ \chi^{\prime}=\frac{\Pi}{H},\\ \\ H^{2}=\frac{\Pi^{2}+2\mathcal{U}}{6M_{P}^{2}},\end{array}\right.\end{array}$ (14)
@@ IV Facing observational bounds
where $\phi$ is the (canonically normalized) Higgs field in the unitary gauge and $\xi$ is a dimensionless coefficient. This is true even for values of $\xi$ that are not large, guaranteeing perturbative unitarity up to scales close to $M_{P}$, where an ultraviolet completion of the effective field theory we are considering should be unavoidably taken into account. Indeed, the term (15) allows for the decay of the pseudoscalaron into two Higgs bosons after inflation, leading to a reheating temperature $T_{\rm RH}\gtrsim 10^{9}|\xi|$ GeV (in the minimal setup considered here) [14].
@@ V Conclusions
According to the new measurements and constraints provided by ATC [1, 2] the class of observationally favored inflationary models has significantly changed compared to the previous Planck and BK18 determinations [3, 4].
In this work it has been shown that a simple explanation of the ATC determinations may be provided if the inflaton is a component of a dynamical connection that is independent of the metric. In this way not only gravity but also inflation has a geometrical origin. This conclusion has been reached by performing a detailed analysis of the inflationary predictions of the minimal model of this sort, whose inflationary action appears in Eq. (1), including $n_{s}$, $r$, $P_{R}$ and $\alpha_{s}$, and by comparing them with the observational constraints.

### v2

@@
Abstract
The Atacama Cosmology Telescope (ACT) has recently released new measurements and constraints on inflationary observables. In this paper it is shown that a component of a dynamical affine connection, which is independent of the metric, can easily drive inflation in agreement with these observations. Such geometrical explanation of inflation is analysed in detail here in the minimal model, including the predictions for the scalar spectral index $n_{s}$ and its running $\alpha_{s}$, the amplitude of the scalar perturbations and the tensor-to-scalar ratio $r$. Furthermore, this minimal model is shown to provide an inflationary attractor: arbitrary initial values of the kinetic energy density are dynamically attracted down to negligible values compared to the potential energy density in homogeneous and isotropic metrics. Also, the role of the Higgs boson during and after inflation is briefly discussed.
@@ I Introduction
In Refs. [1, 2] the ACT collaboration has provided new determinations of several inflationary observables. One of the most evident differences with respect to the values previously furnished by the Planck and BICEP/Keck (BK18) collaborations [3, 4] is a higher favored value for $n_{s}$: when the new ACT observations are combined with Planck data and baryon acoustic oscillation data from the Dark Energy Spectroscopic Instrument (DESI) [5] (this combination is denoted P-ACT-LB) Ref. [1] finds $n_{s}=0.9743\pm 0.0034$.
@@
The main purpose of this paper is to offer a well-motivated scenario that can explain the new results of Refs. [1, 2]. While Einstein’s general relativity (GR) explains gravity in geometrical terms (the distances are measured through the metric and the gravitational force is determined by the affine connection), inflation is usually accounted for by adding an ad hoc scalar field, which is unrelated to geometry. However, from the purely geometrical point of view the metric and the connection, unlike in GR, can be completely independent objects and, moreover, can contain extra degrees of freedom besides the spin-2 graviton. This generalized scenario is known as metric-affine gravity (see [12, 13] for detailed discussions and further references). The extra degrees of freedom carried by the affine connection can include the inflaton [14, 15, 16] and this provides a geometrical origin also for this scalar field and the subsequent period of reheating too [14, 15], which sets the initial conditions for the radiation dominated phase of the universe.
Addressing the compatibility of such scenario with the new ACT results is not the only purpose of this work. Other new goals include studying the implications of the couplings between the inflaton and the Standard Model particles for reheating and inflation, investigating the presence of an inflationary attractor and studying the running of the spectral index in this scenario.
@@ II The model
[…]
@@
[…]
@@ III Inflationary predictions
$N(\omega)=\frac{1}{M_{P}^{2}}\int^{\omega}d\omega^{\prime}\,\mathcal{U}\left(\frac{d\mathcal{U}}{d\omega^{\prime}}\right)^{-1}$ (9)
@@
$\xi_{2}(\omega)=\frac{64M_{P}^{4}\cosh^{2}(X(\omega))\left(M_{P}^{2}\sinh(X(\omega))-\beta\right)}{9\left(M_{P}^{2}\sinh(X(\omega))-4\beta\right)^{3}}.\\$ (13)
@@
[…]
@@ IV Facing observational bounds
In Fig. 3  $\mathcal{U}$ is plotted as a function of $\omega$ for $\beta=-20M_{P}^{2}$. One can note that the values of the energy density, $\sim\mathcal{U}$, are well below the Planck scale, the maximal cutoff of the theory (as Einstein’s gravity breaks down at that scale). Therefore, if the UV completion of the present model occurs at energies not much smaller than $M_{P}$, the effective field theory approach used here is expected to provide a good approximation for the observable quantities.
Figure 3: The potential density as a function of the canonically normalized field. Here $\beta=-20M_{P}^{2}$. For all curves $c$ is fixed to reproduce $P_{R}=2.1\times 10^{-9}$. Also, the black points correspond to the values of the inflaton for which $N_{e}=49$ (upper curve) and $N_{e}=60$ (lower curve); the corresponding predictions for $n_{s}$ and $r$ are also shown.
@@
where $\phi$ is the (canonically normalized) Higgs field in the unitary gauge and $\xi$ is a dimensionless coefficient. This is true even for values of $\xi$ that are not large, guaranteeing perturbative unitarity up to scales close to $M_{P}$, where an ultraviolet completion of the effective field theory we are considering should be unavoidably taken into account. Indeed, the term (15) allows for the decay of the pseudoscalaron into two Higgs bosons after inflation, leading to a reheating temperature $T_{\rm RH}\gtrsim 10^{9}|\xi|$ GeV (in the minimal setup considered here and for $|\beta|\gtrsim M_{P}^{2}$[14]. Note that for $|\xi|\gtrsim 1$ the reheating temperature is very large, $T_{\rm RH}\gtrsim 10^{9}$ GeV, and even for a very small $|\xi|$ the reheating temperature remains much larger than what is required by observations.
@@ V Conclusions
According to the new measurements and constraints provided by ACT [1, 2] the class of observationally favored inflationary models has significantly changed compared to the previous Planck and BK18 determinations [3, 4].
In this work it has been shown that a simple explanation of the ACT determinations may be provided if the inflaton is a component of a dynamical connection that is independent of the metric. In this way not only gravity but also inflation has a geometrical origin. This conclusion has been reached by performing a detailed analysis of the inflationary predictions of the minimal model of this sort, whose inflationary action appears in Eq. (1), including $n_{s}$, $r$, $P_{R}$ and $\alpha_{s}$, and by comparing them with the observational constraints.
@@
The values of $\alpha_{s}$ predicted by the model are very small: as illustrated in Fig. 4, any evidence for $\alpha_{s}$ significantly above the $10^{-4}$ order of magnitude would falsify the model.
Interestingly, the requirement that the two scales of the model with mass dimension, $M_{P}$ and $|\sqrt{\beta}|$, be close to each other (which appears natural) favours values of $n_{s}$ close to the upper observational bound. This can be tested by future experiments such as LiteBIRD [35].